OpenStudy (anonymous):

why does: y+dy=(x+dx)^-2 equal: =x^-2(1+dx/x)^-2

OpenStudy (anonymous):

put x in evidence

OpenStudy (atlas):

Another way of thinking is divide the expression on R.H.S with x^-2 and then multiply it with x^-2 on the outside

OpenStudy (atlas):

multiplying and dividing Right hand side with x^-2 \[\frac{ x ^{-2}(x+dx)^{-2} }{ x^{-2} }\] \[x^{-2} (\frac{ x+dx }{ x })^{-2}\]

OpenStudy (atlas):

simplify it to x^-2 (1+dx/x)^-2 I hope it makes sense now

OpenStudy (anonymous):

I stiil dont get it Could someone please explain it step by step?

myininaya (myininaya):

Ok, do you agree that x+dx could be written as x(1+dx/x)?

myininaya (myininaya):

\[x+dx=x+\frac{x}{x} dx=x+x \cdot \frac{dx}{x}=x(1+\frac{dx}{x})\]

OpenStudy (anonymous):

how can x+dx be written as x(1+dx/x)?

myininaya (myininaya):

\[(x+dx)^{-2}=[(x)(1+\frac{dx}{x})]^{-2}=x^{-2}(1+\frac{dx}{x})^{-2} \]

myininaya (myininaya):

Did you see what I said above what you said?

myininaya (myininaya):



myininaya (myininaya):